OpenStudy (anonymous):

Find the indicated derivative. In this case, the independent variable is a (unspecified) function of t. V = 8/7 (3.14r^2)

OpenStudy (anonymous):

is this the original function? \[V = \frac{8}{7}\pi r^2\]

OpenStudy (anonymous):

@Mathmuse yes

OpenStudy (anonymous):

actually its squared by 7, not 2

zepdrix (zepdrix):

\[\large V=\frac{ 8 }{ 7 }\pi r^7\] Like that? :o

OpenStudy (anonymous):

@zepdrix yes

zepdrix (zepdrix):